OpenStudy (anonymous):

. On a map, 1/3 inch equals 15 miles. The distance between two towns on a map is 3 2/3 inches. How many miles are actually between the two towns? i tried making this into a proportion .33in/15m= 3.67in/x then cross multiplied .33x=5505 now im stuck bc this is kinda weird lol! help! i know the answer which is 165 but i dont know how to work the problem.

OpenStudy (anonymous):

do 1 inch equals 45 miles so 45 * 3 = x then 15 * 2 (for the 2/3) = y then x + y

OpenStudy (anonymous):

why 1 inch

OpenStudy (anonymous):

where did u get 45? whhhattttt

OpenStudy (anonymous):

1/3 of an inch = 15 miles right? so 3 * 15 = 45 miles per inch so 45 * 3 for the 3 inches then the 2/3 inches is 30 miles so 45*3 + 30

OpenStudy (anonymous):

can u not work this problem in a porpotion?

OpenStudy (anonymous):

umm idk, i find my way to be easier if u have to set it up as a proportion im sure u could

OpenStudy (anonymous):

well thats how i wanna do it bc im comfortable with proportions

OpenStudy (anonymous):

its the same thing lmao

OpenStudy (anonymous):

no nu uh not in my brain

OpenStudy (anonymous):

ok so draw it out on the thingy if its the same thing!

OpenStudy (anonymous):

see at the bottom where it says draw

OpenStudy (anonymous):

Alright, Since there are Three 1/3's in a whole 1. And 1/3 is 15 Miles, You would know that you need to multiply the 15 by 3, since its easier to deal with a whole number. Understand that part?

OpenStudy (anonymous):

kyle, uhm not exactly lol im slow so sorry

OpenStudy (anonymous):

reality vs perception... regardless of if its not the same in ur brain its the same in the majority of people thus thats the reality of it because reality is a mass collection of perception, either way 1/3 = 15 so 3 2/3 = x

OpenStudy (anonymous):

i want to work this in a proportion tho

OpenStudy (anonymous):